# Enhancing Semantic Segmentation with Continual Self-Supervised Pre-training

### v1

@@ Abstract
Self-supervised learning (SSL) has emerged as a central paradigm for training foundation models by leveraging large-scale unlabeled datasets, often producing representations with strong generalization capabilities. These models are typically pre-trained on general-purpose datasets such as ImageNet and subsequently adapted to various downstream tasks through finetuning. While recent advances have explored parameter-efficient strategies for adapting pre-trained models, extending SSL pre-training itself to new domains—particularly under limited data regimes and for dense prediction tasks—remains underexplored. In this work, we address the problem of adapting vision foundation models to new domains in an unsupervised and data-efficient manner, specifically targeting downstream semantic segmentation. We propose GLARE (Global Local and Regional Enforcement), a novel continual self-supervised pre-training task designed to enhance downstream segmentation performance. GLARE introduces patch-level augmentations to encourage local consistency and incorporates a regional consistency constraint that leverages spatial semantics in the data. For efficient continual pre-training, we initialize Vision Transformers (ViTs) with weights from existing SSL models and update only lightweight adapter modules—specifically UniAdapter—while keeping the rest of the backbone frozen. Experiments across multiple semantic segmentation benchmarks on different domains demonstrate that GLARE consistently improves downstream performance with minimal computational and parameter overhead.
@@ 1 Introduction
Self-supervised learning (SSL) has revolutionized the training of foundation models by enabling the extraction of rich, generalizable features from vast unlabeled datasets (Caron et al., 2021; Chen et al., 2021; Oquab et al., 2024; Assran et al., 2023). This paradigm has proven to be particularly valuable in computer vision, where the abundance of unlabeled images can be leveraged to learn robust visual representations without the need for expensive manual annotations. Recent advances in SSL have introduced sophisticated techniques, spanning from innovative data augmentation strategies (Grill et al., 2020) and masked image modeling (He et al., 2022) to enhanced feature matching through self-attention mechanisms (Su et al., 2023; Su & Ji, 2024) and refined loss functions.
While these developments have led to powerful models pre-trained on extensive generic datasets like ImageNet (Deng et al., 2009), they often fall short when confronted with specialized technical domains. This limitation becomes particularly apparent in real-world scenarios where domain-specific data is scarce, especially labeled data. The challenge is further compounded when considering dense prediction tasks like semantic segmentation, which demand more fine-grained semantic understanding compared to classification tasks. These issues have therefore sparked significant interest in adaptation techniques, particularly finetuning and continual learning to bridge the gap between generic pre-training and domain-specific applications.
Our research focuses on enhancing foundation model features for task-specific limited data scenarios through continual pre-training within a pure SSL framework. While previous research has explored this direction by training batch normalization layers using conventional SSL approaches for classification tasks (Reed et al., 2022), these studies do not fully investigate other aspects of the SSL paradigm, such as augmentations and matching, in limited data scenarios. Furthermore, while extensive research exists on continual pre-training for classification datasets (Lin et al., 2022; Cheng et al., 2023; Reed et al., 2022; Tang et al., 2024), the distinct challenges posed by semantic segmentation warrant separate investigation, as features of foundation models often perform differently between classification and segmentation tasks (Su & Ji, 2024; Caron et al., 2021; Oquab et al., 2024).
@@
We provide model-specific experiments with existing state-of-the-art foundation models, while comparing our proposed SSL with standard SSL approaches. We validate our approach through experiments across four datasets: ADE20k (Zhou et al., 2017), Pascal Context (Mottaghi et al., 2014), Cityscapes (Cordts et al., 2016) and LoveDA (Wang et al., 2021a) (satellite images), showing performance improvements on downstream segmentation task after applying the proposed SSL approach in continual pre-training.
@@ 2.1 SSL for global image understanding
SSL methods for vision aim to learn rich visual representations from large-scale unlabeled data by enforcing various learning objectives and paradigms. One prominent paradigm is joint embedding networks, exemplified by methods like MoCo (He et al., 2020) and DINO (Caron et al., 2021). MoCo employs contrastive learning with a memory bank for gathering negative examples (He et al., 2020; Wu et al., 2018), while later approaches, such as SimCLR (Chen et al., 2020b) and MoCo-V3 (Chen et al., 2021), eliminate the memory bank and use training batch samples as negatives. In contrast, methods like DINO (Caron et al., 2021), BYOL (Grill et al., 2020), and SimSiam (Chen & He, 2021) forego negative pairs altogether, focusing instead on positive pair similarity. Techniques for achieving this include contrastive learning (e.g., ZeroCL (Zhang et al., 2022)), clustering (e.g., SwAV (Caron et al., 2020), SeLa (Asano et al., 2020), MSN (Assran et al., 2022)), and alternative objectives such as redundancy reduction in Barlow Twins (Zbontar et al., 2021).
Another paradigm, inspired by Natural Language Processing (NLP) (Radford & Narasimhan, 2018; Devlin et al., 2019), is generative SSL. Examples include iGPT (Chen et al., 2020a), which reconstructs masked pixels, and vision-specific approaches like BEiT (Bao et al., 2022) and MAE (He et al., 2022), which reconstruct masked patches. More recently, I-JePa (Assran et al., 2023) advanced this concept by predicting embeddings of masked patches using contextual and target encoders. Hybrid approaches, such as iBoT (Zhou et al., 2022), DINOv2 (Oquab et al., 2024), combine joint embedding learning with masked image modeling techniques like those in MAE (He et al., 2022).
@@ 2.2 SSL for dense prediction
SSL methods tailored at dense prediction tasks, including semantic segmentation, have gained special attention in recent years. These methods can be grouped into three main categories:
@@
This group of methods focuses on similarity within regions or at the object level. ReSim (Xiao et al., 2021) aligns representations of overlapping sliding window regions across views. At the object level, methods like SoCo (Wei et al., 2021), ORL (Xie et al., 2021b), and SCRL (Roh et al., 2021) employ techniques such as selective search which are very expensive. SelfPatch (Yun et al., 2022) defines regions as the set of patches in the direct neighborhood of a central patch and enforces similarity between these neighbors. FLSL (Su et al., 2023) on the other hand introduces intra- and inter-view clustering, attracting representations of the same concept while repelling clusters of different concepts across augmentations. Finally, the most recent work UDI (Su & Ji, 2024) encourages multimodal local predicitions by adding an additional class token and solve the semantic misalignment problem.
In this work, we use UDI pre-trained model to initialize our backbones for continual pre-training as UDI (Su & Ji, 2024) showcases strong results in SSL for dense prediction tasks. We then leverage the best performing model and show consistent improvements using our pre-training strategy in continual pre-training for semantic segmentation.
@@ 2.3 Continual pre-training
Recent works have explored unsupervised continual pre-training as a means to adapt large-scale pre-trained models to new domains. These approaches aim to bridge the gap between general-purpose pre-training and downstream domain-specific tasks by refining representations learned on source domains. Hierarchical Pre-Training (HPT) (Reed et al., 2022) introduces a framework for self-supervised continual pre-training, where a model trained on source domain data is further pre-trained on target domain data by finetuning all model weights in a sequential framework. Other approaches adapts ViTs using masked image modeling on target domain (Mendieta et al., 2023). More recent efforts have explored parameter-efficient adaptation strategies, such as incorporating lightweight modules like adapters (e.g., LoRA) to reduce compute cost (Scheibenreif et al., 2024; Khanna et al., 2025). While these methods demonstrate the promise of continual pre-training, they are primarily developed for downstream classification tasks and rely on large-scale datasets (typically exceeding 100k images). In contrast, our work focuses on designing a continual self-supervised pre-training pipeline tailored to dense prediction tasks, specifically semantic segmentation, in data-scarce target domains.
@@ 3.1 Vision Transformers
Let $X\in\mathbb{R}^{C\times H\times W}$ be an image, where $H$ and $W$ represent the height and width of the image, respectively, and $C$ the number of channels. A Vision Transformer model (ViT) (Dosovitskiy et al., 2020) considers the image $x$ as a set of $N$ non overlapping patches $x_{i}\in\mathbb{R}^{CP^{2}}$, of resolution $P\times P$ and with $C$ channels. These patches are then projected through a linear layer to a space of dimension $D$, such that $z_{i}=Wx^{(i)}+E^{i}_{pos}$, where $W\in\mathbb{R}^{D\times CP^{2}}$ is a linear projection and $E^{i}_{pos}\in\mathbb{R}^{D}$ is the positional embedding for the patch at index $i$. A learnable token $z_{[CLS]}\in\mathbb{R}^{D}$, referenced as [CLS], is prepended to the sequence of patches to extract global information from the image. The resulting input sequence is thus defined as $z=[z_{[CLS]},z_{1},z_{2},\ldots,z_{N}]$. Then, ViTs take the input $e$ to produce global level ($e_{[CLS]}$) and patch level ($e_{i}$) representations by using its encoder. In the same way as in (Yun et al., 2022; Zhang et al., 2023b), we refer to the encoder as $f_{\theta}$ with parameters $\theta$ and we use Equation˜2 to represent the whole process of a ViT:
@@ 3.2 Continual pre-training with adapters
Our interest is to improve the SSL ViT features via a task-agnostic SSL framework in the continual learning setup. Therefore, we limit our work’s investigation to finding the right SSL framework for continual learning rather than exploring different continual learning paradigms. Among the various continual learning approaches, adapter-based (Ding et al., 2022; Lu et al., 2023; Hu et al., 2022), and memory or replay based (Winter et al., 2024; Reed et al., 2022; Wang et al., 2024) strategies are considered to be state-of-the-art. However, due to their simplicity and compatibility with a wide range of architectures, adapter-based methods are widely used in continual learning (Scheibenreif et al., 2024; Khanna et al., 2025) and finetuning. We therefore consider a simple adapter named UniAdapter (Lu et al., 2023) composed of two linear layers with activation for adapting features after every self-attention layer in ViT. Therefore, given a pre-trained SSL model, our investigation involves training only the adapter layers while freezing the model weights, following the standard practice. Note that, by using the parameter-efficient adapter used typically for finetuning models, our problem formulation departs from previously used SSL continual learning, for which all network layers are trained (Reed et al., 2022). Our choice is mainly motivated by the need to isolate natural catastrophic forgetting from the suitability of the SSL framework (see  Table˜2).
@@ 3.3 Problem formulation
Given an encoder trained via SSL (Su & Ji, 2024; Caron et al., 2021), we are interested in improving the output feature embedding $e$ by training only the adapter parameters $\theta_{A}$ via SSL. Thus we write the new ViT parameters as $\theta_{A}$, resulting in a different feature embedding $e^{\prime}$ given an input image. Here $\theta_{A}$ denotes the trainable parameters of the $\operatorname{Adapter}$ layers. Given a new target dataset $\mathcal{T}$, our problem is to study and develop the most suitable SSL framework that maximizes the performance of $e^{\prime}$. We measure the performance of the feature embedding $e^{\prime}$ by finetuning our model for semantic segmentation together with an additional decoder. In other words, we use semantic segmentation finetuning as the representative metric for gauging feature embeddings $e^{\prime}$.
4 Method: GLARE pre-training
@@
GLARE (Global Local and Regional Enforcement) is an SSL framework focused on learning representations at different levels during the pre-training process. When we as humans look at a picture, in practice we focus our attention to varying levels of detail (Navon, 1977; Shi et al., 2014). We can summarize these levels as 1) the image as a whole, 2) as a set of regions/objects, and 3) as a collection of specific details on these regions, which can be encoded into patches or pixels. GLARE, inspired by this concept, is designed as a three-level pre-training strategy, combining the enforcement of global, regional, and local consistency for learning coarse to fine-grained representations, which are crucial to develop a more in-depth understanding of image data. We adopt the usual student-teacher framework for self-supervised learning described in Caron et al. (2021); He et al. (2020); Chen et al. (2021); Yun et al. (2022); Zhang et al. (2023b). Following the same naming convention as in Chen et al. (2021), we consider two ViT encoders: a base encoder $f_{\theta,\theta_{A}}$ and a momentum encoder $f_{\theta^{\prime},\theta^{\prime}_{A}}$, parameterized by ($\theta,\theta_{A}$) and ($\theta^{\prime},\theta^{\prime}_{A}$), respectively. The parameters related to the momentum encoder ($\theta^{\prime},\theta^{\prime}_{A}$) are updated through an exponential moving average (EMA) of those of the base encoder ($\theta,\theta_{A}$), as in He et al. (2020). Figure˜1 describes GLARE pre-training strategy. Specifically, in our continual pre-training setup, we train only the adapter parameters $\theta_{A}$ of the base encoder and update those of the momentum encoder $\theta^{\prime}_{A}$ through EMA as stated in Section˜3.3. In the following sections, we describe more in detail the three levels of consistency enforced.

### v2

@@ Abstract
Self-supervised learning (SSL) has emerged as a central paradigm for training foundation models by leveraging large-scale unlabeled datasets, often producing representations with strong generalization capabilities. These models are typically pre-trained on general-purpose datasets such as ImageNet and subsequently adapted to various downstream tasks through finetuning.
While prior work has investigated parameter-efficient adaptation methods like adapters, LoRA, and prompt tuning, primarily targeting downstream finetuning, extending the SSL pre-training itself in a continual manner to new domains under limited data remains largely underexplored, especially for downstream dense prediction tasks like semantic segmentation. In this work, we address the challenge of adapting vision foundation models to low-data target domains through continual self-supervised pre-training, specifically targeting downstream semantic segmentation. We propose GLARE (Global Local and Regional Enforcement), a novel continual self-supervised pre-training task designed to enhance downstream semantic segmentation performance. GLARE introduces patch-level augmentations to encourage local consistency and incorporates a regional consistency constraint that leverages spatial semantics in the data. For efficient continual pre-training, we initialize Vision Transformers (ViTs) with weights from existing SSL models and update only lightweight adapter modules–specifically UniAdapter–while keeping the rest of the backbone frozen. Experiments across multiple semantic segmentation benchmarks on different domains demonstrate that GLARE consistently improves downstream performance with minimal computational and parameter overhead.
@@ 1 Introduction
Self-supervised learning (SSL) has revolutionized the training of foundation models by enabling the extraction of rich, generalizable features from vast unlabeled datasets (Caron et al., 2021; Chen et al., 2021; Oquab et al., 2024; Assran et al., 2023). This paradigm has proven to be particularly valuable in computer vision, where the abundance of unlabeled images can be leveraged to learn robust visual representations without the need for expensive manual annotations. Recent advances in SSL have introduced sophisticated techniques, spanning from innovative data augmentation strategies (Grill et al., 2020) and masked image modeling (He et al., 2022) to enhanced feature matching through self-attention mechanisms (Su et al., 2023; Su and Ji, 2024) and refined loss functions.
While these developments have led to powerful models pre-trained on extensive generic datasets like ImageNet (Deng et al., 2009), they often fall short when confronted with specialized technical domains. This limitation becomes particularly apparent in real-world scenarios where domain-specific data is scarce, especially labeled data. The challenge is further compounded when considering dense prediction tasks like semantic segmentation, which demand more fine-grained semantic understanding compared to classification tasks. These issues have therefore sparked significant interest in adaptation techniques, particularly finetuning, and continual learning to bridge the gap between generic pre-training and domain-specific applications.
Our research focuses on adapting a generalist SSL pre-trained backbone to a new domain with limited data through continual pre-training within a pure SSL framework. Unlike continual learning, which typically refers to adapting a model trained for a specific downstream task to a new domain or task, our focus is on the pre-training stage: specifically, we start from a model already pre-trained with SSL on some data and we continue the unsupervised SSL pre-training on new, unlabeled domain data. This allows the model to better align its representations with the target domain using the available unlabeled data prior to the downstream adaptation stage, that then makes use of the scarce available labeled data. Previous work has explored aspects of continual pre-training by, for example, training batch normalization layers using conventional SSL approaches for classification tasks (Reed et al., 2022). However, these studies do not fully investigate other key components of the SSL paradigm, such as augmentations and matching, particularly in limited data regimes. Furthermore, while extensive research exists on continual pre-training for image classification (Lin et al., 2022; Cheng et al., 2023; Reed et al., 2022; Tang et al., 2024), the distinct challenges posed by semantic segmentation warrant separate investigation, as features of foundation models often perform differently between classification and segmentation tasks (Su and Ji, 2024; Caron et al., 2021; Oquab et al., 2024).
@@
We provide model-specific experiments with existing state-of-the-art foundation models, while comparing our proposed SSL with standard SSL approaches. We validate our approach through experiments across four datasets: ADE20k (Zhou et al., 2017), Pascal Context (Mottaghi et al., 2014), Cityscapes (Cordts et al., 2016) and LoveDA (Wang et al., 2021a) (satellite images), showing consistent performance improvements on downstream segmentation task after applying the proposed SSL approach in continual pre-training.
@@ 2.1 SSL for global image understanding
SSL methods for vision aim to learn rich visual representations from large-scale unlabeled data by enforcing various learning objectives and paradigms. One prominent paradigm is joint embedding networks, exemplified by methods like MoCo (He et al., 2020) and DINO (Caron et al., 2021). MoCo employs contrastive learning with a memory bank for gathering negative examples (He et al., 2020; Wu et al., 2018), while later approaches, such as SimCLR (Chen et al., 2020b) and MoCo-V3 (Chen et al., 2021), eliminate the memory bank and use training batch samples as negatives. In contrast, methods like DINO (Caron et al., 2021), BYOL (Grill et al., 2020), and SimSiam (Chen and He, 2021) forego negative pairs altogether, focusing instead on positive pair similarity. Techniques for achieving this include contrastive learning (e.g., ZeroCL (Zhang et al., 2022)), clustering (e.g., SwAV (Caron et al., 2020), SeLa (Asano et al., 2020), MSN (Assran et al., 2022)), and alternative objectives such as redundancy reduction in Barlow Twins (Zbontar et al., 2021).
Another paradigm, inspired by Natural Language Processing (NLP) (Radford et al., 2018; Devlin et al., 2019), is generative SSL. Examples include iGPT (Chen et al., 2020a), which reconstructs masked pixels, and vision-specific approaches like BEiT (Bao et al., 2022) and MAE (He et al., 2022), which reconstruct masked patches. More recently, I-JePa (Assran et al., 2023) advanced this concept by predicting embeddings of masked patches using contextual and target encoders. Hybrid approaches, such as iBoT (Zhou et al., 2022), DINOv2 (Oquab et al., 2024), combine joint embedding learning with masked image modeling techniques like those in MAE (He et al., 2022).
@@ 2.2 SSL for dense prediction
SSL methods tailored to dense prediction tasks, including semantic segmentation, have gained special attention in recent years. These methods can be grouped into three main categories:
@@
This group of methods focuses on similarity within regions or at the object level. ReSim (Xiao et al., 2021) aligns representations of overlapping sliding window regions across views. At the object level, methods like SoCo (Wei et al., 2021), ORL (Xie et al., 2021b), and SCRL (Roh et al., 2021) employ techniques such as selective search which are computationally expensive. SelfPatch (Yun et al., 2022) defines regions as the set of patches in the direct neighborhood of a central patch and enforces similarity between these neighbors. FLSL (Su et al., 2023) on the other hand, introduces intra- and inter-view clustering, attracting representations of the same concept while repelling clusters of different concepts across augmentations. Finally, the most recent work UDI (Su and Ji, 2024) encourages multimodal local predicitions by adding an additional class token and solve the semantic misalignment problem.
In this work, we use UDI pre-trained model to initialize our backbones for continual pre-training as UDI (Su and Ji, 2024) showcases strong results in SSL for dense prediction tasks. We then leverage the best performing model and show consistent improvements using our pre-training strategy in continual pre-training for semantic segmentation.
@@ 2.3 Continual pre-training
Recent works have explored unsupervised continual pre-training as a means to adapt large-scale pre-trained models to new domains. These approaches aim to bridge the gap between general-purpose pre-training and downstream domain-specific tasks by refining representations learned on source domains. Hierarchical Pre-Training (HPT) (Reed et al., 2022) introduces a framework for self-supervised continual pre-training, where a model trained on source domain data is further pre-trained on target domain data by finetuning all model weights in a sequential framework. Other approaches adapts ViTs using masked image modeling on target domain (Mendieta et al., 2023). More recent efforts have explored parameter-efficient adaptation strategies, such as incorporating lightweight modules like adapters (e.g., LoRA) to reduce compute cost (Scheibenreif et al., 2024; Khanna et al., 2024). While these methods demonstrate the promise of continual pre-training, they are primarily developed for downstream classification tasks and rely on large-scale datasets (typically exceeding 100k images). In contrast, our work focuses on designing a continual self-supervised pre-training pipeline tailored to dense prediction tasks, specifically semantic segmentation, in data-scarce target domains. Rather than proposing a new parameter-efficient tuning method, we repurpose an adapter-based approach and apply it during the continual self-supervised pre-training stage, enabling domain-aligned representation learning prior to downstream adaptation.
@@ 3.1 Vision Transformers
Let $X\in\mathbb{R}^{C\times H\times W}$ be an image, where $H$ and $W$ represent the height and width of the image, respectively, and $C$ the number of channels. A Vision Transformer model (ViT) (Dosovitskiy et al., 2021) considers the image $x$ as a set of $N$ non overlapping patches $x_{i}\in\mathbb{R}^{CP^{2}}$, of resolution $P\times P$ and with $C$ channels. These patches are then projected through a linear layer to a space of dimension $D$, such that $z_{i}=Wx^{(i)}+E^{i}_{pos}$, where $W\in\mathbb{R}^{D\times CP^{2}}$ is a linear projection and $E^{i}_{pos}\in\mathbb{R}^{D}$ is the positional embedding for the patch at index $i$. A learnable token $z_{[CLS]}\in\mathbb{R}^{D}$, referenced as [CLS], is prepended to the sequence of patches to extract global information from the image. The resulting input sequence is thus defined as $z=[z_{[CLS]},z_{1},z_{2},\ldots,z_{N}]$. Then, ViTs take the input to produce global level ($e_{[CLS]}$) and patch level ($e_{i}$) representations by using its encoder. In the same way as in (Yun et al., 2022; Zhang et al., 2023b), we refer to the encoder as $f_{\theta}$ with parameters $\theta$ and we use Equation˜2 to represent the whole process of a ViT:
@@ 3.2 Continual pre-training with adapters
Our interest is to improve the SSL ViT features via a task-agnostic SSL framework in the continual learning setup. Therefore, we limit our work’s investigation to finding the right SSL framework for continual learning rather than exploring different continual learning paradigms. Among the various continual learning approaches, adapter-based (Ding et al., 2022; Lu et al., 2023; Hu et al., 2022), and memory or replay based (Winter and others, 2024; Reed et al., 2022; Wang et al., 2024) strategies are considered to be state-of-the-art. However, due to their simplicity and compatibility with a wide range of architectures, adapter-based methods are widely used in continual learning (Scheibenreif et al., 2024; Khanna et al., 2024) and finetuning. We therefore consider a simple adapter named UniAdapter (Lu et al., 2023) composed of two linear layers with activation for adapting features after every self-attention layer in ViT. Therefore, given a pre-trained SSL model, our investigation involves training only the adapter layers while freezing the model weights, following the standard practice. Note that, by using the parameter-efficient adapter used typically for finetuning models, our problem formulation departs from previously used SSL continual learning, for which all network layers are trained (Reed et al., 2022). Our choice is mainly motivated by the need to isolate natural catastrophic forgetting from the suitability of the SSL framework (see  Table˜2).
@@ 3.3 Problem formulation
Given an encoder trained via SSL (Su and Ji, 2024; Caron et al., 2021), we are interested in improving the output feature embedding by training only the adapter parameters $\theta_{A}$ via SSL. Given an input image, we denote with $e^{\prime}$ the output feature embedding of the new ViT with the Adapter. Here $\theta_{A}$ denotes the trainable parameters of the $\operatorname{Adapter}$ layers. Given a new target dataset $\mathcal{T}$, our problem is to study and develop the most suitable SSL framework that maximizes the performance of $e^{\prime}$. We measure the performance of the feature embedding $e^{\prime}$ by finetuning our model for semantic segmentation together with an additional decoder. In other words, we use semantic segmentation finetuning as the representative metric for gauging the feature embeddings $e^{\prime}$. We highlight that this setting differs from traditional supervised continual learning. Here, the continual aspect refers to the sequential refinement of the pre-trained weights $\theta_{A}$ of the ViT backbone on a new unlabeled distribution $\mathcal{T}$.
4 Method: GLARE Continual Pre-Training
@@
GLARE (Global Local and Regional Enforcement) is an SSL framework focused on learning representations at different levels during the pre-training process. When we as humans look at a picture, in practice we focus our attention to varying levels of detail (Navon, 1977; Shi et al., 2014). We can summarize these levels as 1) the image as a whole, 2) as a set of regions/objects, and 3) as a collection of specific details on these regions, which can be encoded into patches or pixels.
GLARE, inspired by this concept, is designed as a three-level pre-training strategy, combining the enforcement of global, regional, and local consistency for learning coarse to fine-grained representations, which are crucial to develop a more in-depth understanding of image data. We adopt the usual student-teacher framework for self-supervised learning described in Caron et al. (2021); He et al. (2020); Chen et al. (2021); Yun et al. (2022); Zhang et al. (2023b). Following the same naming convention as in Chen et al. (2021), we consider two ViT encoders: a base encoder $f_{\theta,\theta_{A}}$ and a momentum encoder $f_{\theta^{\prime},\theta^{\prime}_{A}}$, parameterized by ($\theta,\theta_{A}$) and ($\theta^{\prime},\theta^{\prime}_{A}$), respectively. The parameters related to the momentum encoder ($\theta^{\prime},\theta^{\prime}_{A}$) are updated through an exponential moving average (EMA) of those of the base encoder ($\theta,\theta_{A}$), as in He et al. (2020). Figure˜1 describes GLARE pre-training strategy. Specifically, in our continual pre-training setup, we train only the adapter parameters $\theta_{A}$ of the base encoder and update those of the momentum encoder $\theta^{\prime}_{A}$ through EMA as stated in Section˜3.3. In the following sections, we describe more in detail the three levels of consistency enforced.